### v1

# Enhancing Agent Learning through World Dynamics Modeling

Abstract
While large language models (LLMs) have been increasingly deployed across tasks in language understanding and interactive decision-making, their impressive performance is largely due to the comprehensive and in-depth domain knowledge embedded within them. However, the extent of this knowledge can vary across different domains. Existing methods often assume that LLMs already possess such comprehensive and in-depth knowledge of their environment, overlooking potential gaps in their understanding of actual world dynamics. To address this gap, we introduce Discover, Verify, and Evolve (DiVE), a framework that discovers world dynamics from a small number of demonstrations, verifies the correctness of these dynamics, and evolves new, advanced dynamics tailored to the current situation. Through extensive evaluations, we analyze the impact of each component on performance and compare the automatically generated dynamics from DiVE with human-annotated world dynamics. Our results demonstrate that LLMs guided by DiVE can make better decisions, achieving rewards comparable to human players in the Crafter environment.
Enhancing Agent Learning through World Dynamics Modeling
Zhiyuan Sun1thanks:   Equal contribution., Haochen Shi [^1] footnotemark: 1, Marc-Alexandre Côté2, Glen Berseth1, Xingdi Yuan2, Bang Liu1 1 Université de Montréal & Mila, Montréal, Canada 2 Microsoft Research, Montréal, Canada
1 Introduction
By absorbing internet-scale knowledge autoregressively, Large Language Models (LLMs) develop a nuanced understanding of world dynamics Achiam et al. (2023); Team et al. (2023); Brown et al. (2020); Touvron et al. (2023). This understanding enables them to perform well on tasks like question answering Yang et al. (2018); Clark et al. (2019); Lin et al. (2021), planning Song et al. (2023); Brown et al. (2020); Huang et al. (2022), and reasoning over commonsense knowledge Yao et al. (2022, 2024); Besta et al. (2024). However, such understanding requires the training dataset to be diverse, precise, and in-depth enough to cover the domain information needed; otherwise, this may result in a knowledge gap between LLMs and the target domain, as illustrated in Figure 1.
As LLMs tend to follow the most common patterns within the dataset Gunasekar et al. (2023), this mandates that the necessary domain-specific information not only appears on the internet but also appears frequently. For downstream tasks that are either newly established or less popular, such a requirement is often infeasible and leads to suboptimal decision-making for LLMs.
Figure 1: The knowledge gap between LLMs and the domain. Although LLMs are trained on internet-level knowledge, they may still lack the necessary domain-specific knowledge.
To make optimal decisions, LLMs require high-quality, domain-specific information, which is often impractical to source and verify within the vast expanse of internet data Dziri et al. (2022a). Evaluating such information necessitates locating it within the dataset and verifying it against ground-truth domain knowledge, a process that is both labor-intensive and error-prone. Consequently, ensuring LLMs possess the necessary comprehensive and accurate knowledge for decision-making remains a formidable challenge.
Even if LLMs have a general understanding of a downstream domain, achieving optimal decision-making requires in-depth knowledge specific to each state. By providing LLMs with such tailored and in-depth information, their understanding of both the environment and the current state can be further improved. For example, mastering the game of Go involves not only understanding the basic rules but also strategies tailored to specific board states. Moreover, this specialized knowledge can vary not only across domains but also within different states of the same domain, making it impractical to collect in an offline manner.
To address these challenges, we propose Discover, Verify, and Evolve (DiVE). Leveraging the concept of World Model Ha and Schmidhuber (2018); Hafner et al. (2020, 2023), DiVE not only discovers and verifies the primitive world dynamics from the demonstrations but also evolves state-specific knowledge for the downstream domain. By providing LLMs with all these dynamics and knowledge, DiVE can bridge the understanding gap between LLMs and the downstream domain, thereby enhancing LLMs’ decision-making abilities.
DiVE comprises three distinct components: i) The Discoverer: This component iteratively uncovers the dynamics of the environment based on the provided demonstrations. It begins by discovering the primitive dynamics of actions and objects, progresses to subtask steps, and finally determines the topological order of subgoals. This curriculum learning method ensures both learning efficiency and a comprehensive understanding of the environment’s dynamics. ii) The Verifier: This component filters out inaccuracies caused by the tendency of LLMs to hallucinate, ensuring that only precise and reliable information is retained. This step is crucial for maintaining the integrity of the knowledge base. iii) The Evolver: This component reasons in-depth, state-appropriate dynamics, such as strategic plays or contextual decisions. It ensures that the agent is equipped with actionable knowledge tailored to the current state, enhancing real-time decision-making.
Within the Crafter environment Hafner (2021), DiVE learns comprehensive and accurate dynamics from demonstrations and guides the agent through the decision-making process by developing in-depth knowledge. This guidance enables the agent to outperform strong baselines, including the best reported systems, and achieve rewards comparable to those of human players. We also provide both quantitative and qualitative analyses that help understanding DiVE’s behavior, demonstrating the effectiveness of the proposed framework.
In summary, our primary contribution is a general framework that autonomously learns world dynamics from demonstrations and guides LLMs in decision-making by evolving contextual knowledge. This approach not only enhances the capabilities of LLMs in specific domains but also provides a versatile foundation for creating agents capable of mastering complex tasks across various environments through continuous learning and refinement.
2 Problem Formulation
Figure 2: Overall pipeline of DiVE. Left: Learning basic game dynamics from offline demonstrations (Section 3.1). We want to highlight the incorrect game dynamics being identified by the Verifier (labeled by $\times$), they are evidence of the LLM hallucinates false facts perhaps because of memorizing Minecraft data. Right: Learning situational strategies from online interactions (Section 3.2). For simplicity, we omit the verbalization process in the right figure.
Environment with Hierarchical Achievements
Each environment with hierarchical achievements, $C_{i}\in\mathbbm{C}$, operates interactively. At every time step $t$, an agent observes $o_{t}\in\mathcal{O}$ and is expected to respond by selecting an action $a_{t}\in\mathcal{A}$. Subsequently, the environment returns a new observation $o_{t+1}$ of the resulting state. Within each $C_{i}$, the achievement graph $\mathcal{G}=(V,E)$ is structured as a directed acyclic graph (DAG). In this graph, each vertex $v\in V$ denotes an achievement and each edge $(u,v)\in E$ indicates that the achievement $v$ has a dependency on the achievement $u$. To unlock achievement $v$, one must first unlock all its prerequisite ancestors. The main objective is to maximize the number of achievements unlocked by satisfying these dependency conditions.
The Crafter Environment
We primarily employ the Crafter environment as a benchmark to evaluate an agent’s performance in a setting characterized by hierarchical achievements. The Crafter environment  Hafner (2021) is a long-horizon environment typically necessitating hundreds of steps to successfully collect a diamond. Within this environment, the agent must not only determine the topological order of the technology tree as illustrated in Appendix E, but also survive by overcoming dangerous creatures and managing diminishing resources.
Challenge: The Knowledge Gap
In this work, we assume a setting where a pre-trained LLM-based agent is used to solve tasks in the Crafter environment. Conceptually, we use $\mathcal{K}_{\text{data}}$ to denote the set of knowledge being embedded in the LLM through its training process (including both pre-training and fine-tuning). We use $\mathcal{K}_{\text{target}}$ to denote the universal set of knowledge required to solve the target domain (i.e., Crafter). For the LLM to be useful, we hope $\mathcal{K}_{\text{rel}}$, the subset of knowledge in $\mathcal{K}_{\text{data}}$ that is relevant to $\mathcal{K}_{\text{target}}$ to be as large as possible. We also hope that $\mathcal{K}_{\text{rel}}$ contains more accurate knowledge $\mathcal{K}_{+}$ than inaccurate knowledge $\mathcal{K}_{-}$, where $\mathcal{K}_{\text{rel}}=\mathcal{K}_{+}\cup\mathcal{K}_{-}$. Therefore, we define three desirable properties to help conceptualize our system design:
  • Recall $R=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{target}}|}$ measures the extent of knowledge required to solve tasks in the target domain is being covered by the LLM’s knowledge. A low recall typically suggests a large distribution gap between LLM’s training data and the target domain.
  • Precision $P=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{rel}}|}$ measures how accurate the LLM’s knowledge is when grounded in the target domain. An example of inaccurate game dynamics is shown in Figure 2 (left), where the LLM hallucinates that defeating skeleton will drop items, which is not true in Crafter. This may because the LLM has memorized some Minecraft-specific data.
  • Depth $D$ measures the abstraction levels of the knowledge representations. Knowledge can be represented at different abstraction levels, ranging from elementary level game rules to high level strategies.
We want to acknowledge that it is nontrivial to directly quantify the exact knowledge overlap between an LLM and a dataset. We adopt the above concepts from Information Retrieval and related literature to guide our system design. In Section 3, we show mathematically how our system improve on all the three dimensions; in Section 4, we design silver standard on quantifying the knowledge overlap, which provides us evidence that our system can indeed learn useful game dynamics and to some extent fills the knowledge gap.
3 Method
In an ideal world, one could alleviate the aforementioned knowledge gap by fine-tuning the LLM to adapt to the target domain. However, this approach often show less practical due to its reliance on abundant annotated data and significant computational overhead Hu et al. (2021); Zheng et al. (2024); Carta et al. (2023); Ouyang et al. (2022). Our framework, DiVE, is designed to bridge the knowledge gap taking into consideration all the three desirable properties (i.e., recall, precision, and depth), but without the need of collecting extensive data from the target domain. It is a prompt-based method that learns world dynamics $\mathcal{W}$ from the environment.
As illustrated in Figure 2, DiVE is initially bootstrapped from a handful of human demonstration trajectories $H=\{\tau_{i}\}_{i=0}^{N}=\{\{o_{t}^{i},a_{t}^{i},r_{t}^{i}\}_{t=0}^{|\tau_{i}|}\}_{i=0}^{N}$. We verbalize an observation $o_{t}$ into the language space as $\tilde{o}_{t}$ using a Verbalizer (i.e., $\tilde{o}_{t}=\text{Verbalizer}(o_{t})$), resulting in transformed trajectories $\tilde{\tau}_{i}=\{\tilde{o}_{t}^{i},a_{t}^{i},r_{t}^{i}\}_{t=0}^{|\tau_{i}|}$. Next, the Discoverer distills a set of world dynamic candidates $\tilde{\mathcal{W}}=\{\tilde{W}_{+},\tilde{W}_{-}\}$ from human demonstrations $H$, where $\tilde{W}_{+}$ and $\tilde{W}_{-}$ represent the correct and inaccurate world dynamic sets, respectively.
Empirically, we find that it is often inevitable to have $\tilde{W}_{-}$ in $\tilde{\mathcal{W}}$, either due to the backbone LLM’s imperfection in discovering meaningful knowledge from trajectory data, or its tendency of hallucination. Therefore, we use a Verifier to filter out potential invalid and conflicting world dynamic candidates within $\tilde{\mathcal{W}}$, we denote the remaining valid dynamics as $\mathcal{W}$. Next, we use an Evolver, which is designed to derive advanced game-playing strategies $\mathcal{I}$ based on the world dynamics $\mathcal{W}$ and the language description of observation $\tilde{o}_{t}$.
The final decision-making process on primitive actions $a_{t}\in\mathcal{A}$ is hierarchically decomposed as planning tasks on sub-goals $SG$, sub-tasks $ST$, and actions $\mathcal{A}$. The planning procedure is further guided by both $\mathcal{W}$ and $\mathcal{I}$. In cases where $\mathcal{W}\neq\varnothing$, $R$, $P$, and $D$ are guaranteed to increase as formulated below:
[…]
The DiVE framework can be split into two stages: an offline dynamics learning stage and an online strategy learning stage.
3.1 Offline Dynamics Learning
The offline dynamics learning procedure aims at bridging the gap between LLM’s understanding and the basic rules of the Crafter world by learning the world dynamics $\mathcal{W}$ as a prior for further decision-making. Rather than extracting world dynamics from human-authored game manuals or handbooks  Wu et al. (2024b, a), which can be difficult to obtain in many real-world scenarios, we explore learning world dynamics $\mathcal{W}$ from experiences $H$, which is arguably more accessible.
Hierarchical Curriculum Learning
Given the varying complexities in learning the dynamics of different elements in Crafter, we adopt curriculum learning Bengio et al. (2009). We propose a sequential learning approach progressing from simple to more challenging dynamics, facilitating effective learning. Specifically, we propose a method to learn the dynamics of each element within the task decomposition hierarchy, denoted as $TD=\{\mathcal{A}+(O),ST,SG\}$, where $O$ is the set of objects in Crafter. Our approach starts with elements of lower abstraction, such as actions $a\in\mathcal{A}$, and progresses to higher abstraction elements, like sub-goals $sg_{i}\in SG$. The sub-goals sequence $SG=[sg_{1},sg_{2},\ldots]$ is an ordered sequence to unlock achievements in the achievement graph $\mathcal{G}=(V,E)$, with $SG=\text{TopologicalSort}(\mathcal{G})$ and each $sg_{i}$ being a vertex in $V$. We use the Discoverer to extract this order from human demonstrations $H$. The sub-tasks $ST$ are defined as the nodes in the achievement graph $G$, i.e., $ST=V$. Achieving a sub-goal $sg_{i}$ may require completing several sub-tasks multiple times. This approach ensures a logical progression through the tasks, enhancing the learning and application of Crafter’s dynamics.
Discoverer
The Discoverer is designed to identify dynamic candidates $\tilde{\mathcal{W}}$ relating to the elements within the task decomposition hierarchy $TD$. A single dynamics discovering step for an element $E\in TD$ involves following three main steps:
  1. 1.
    Construction of a Semantic Experience Bank ($B^{E}$): For each element $E$, a semantic experience bank $B^{E}$ is constructed from human demonstrations $H$. This bank stores experiences that are transformed from $H$ into a suitable granularity for analyzing dynamics related to $E$. The transformation involves chunking and summarizing the verbalized human demonstrations $\tilde{H}$ to capture essential semantic details.
  2. 2.
    Sampling of Relevant Experiences: For each attribute of an instance $e\in E$, a subset of experiences $B_{e}^{E}$ that are relevant to the instance $e$ is sampled from $B^{E}$.
  3. 3.
    Identification of Dynamic Candidates: A dynamic candidate $\tilde{w}$ is identified from the subset $B_{e}^{E}$ by recognizing patterns that are consistent across all experiences within $B_{e}^{E}$.
The action-level semantic experience bank, denoted as $B^{\mathcal{A}}$, stores transition tuples derived from verbalized human demonstrations $\tilde{H}$, specifically $B^{\mathcal{A}}=\{\{\tilde{o}t,a_{t},\tilde{o}{t+1}\}_{i}\}_{i=1}^{|B^{\mathcal{A}}|}$. Similarly, the object-level semantic experience bank $B^{O}$ collects individual observations related to specific object interactions from $\tilde{H}$, stored as $B^{O}=\{\tilde{o}i\}_{i=1}^{|B^{O}|}$. The sub-task-level semantic experience bank $B^{ST}$ aggregates trajectory segments that represent the completion of sub-tasks, formatted as $B^{ST}=\{\{\tilde{o}_{t},\ldots,a_{tst},\tilde{o}_{tst+1}\}_{i}\}_{i=1}^{|B^{ST}|}$, where $tst$ denotes the time step at which a sub-task $st\in ST$ is completed. For sub-goals, the sub-goal-level experience bank $B^{SG}$ records sequences of sub-tasks that lead to the completion of sub-goals, expressed as $B^{SG}=\{\{st_{t},\ldots,st_{tsg}\}_{i}\}_{i=1}^{|B^{SG}|}$, where $tsg$ is the time step at which the sub-goal $sg$ is achieved.
For action-level dynamics learning, the relevant experiences $B_{a}^{\mathcal{A}}$ are compiled by randomly sampling transition tuples from $B^{\mathcal{A}}$ where the action $a$ is successfully executed. A similar approach is applied to the dynamic learning for other elements within the task decomposition hierarchy $TD$.
For action-level dynamics discovery, we identify the prerequisites and outcomes of each action (e.g., The action MakeWoodPickaxe requires 1 wood). For object-level dynamics, we concentrate on the co-occurrence relationships between objects and their occurrences over time. The attribute set for a sub-task typically includes the general steps required to complete the task, along with its prerequisites, outcomes, and termination conditions. The only attribute of interest for a sub-goal is its correct position within the sub-goal sequence $SG$.
Verifier
Dynamic discovery processes are susceptible to noise arising from several sources, such as confounders, hallucinations by the LLM, and imperfections in the LLM’s ability to derive meaningful knowledge from trajectory data. To mitigate these issues, we introduce a dynamic verifier designed to filter out all noisy dynamic candidates from $\tilde{\mathcal{W}}$. For each dynamic candidate $w\in\tilde{\mathcal{W}}$, the Verifier initiates the process by sampling a subset of relevant semantic experiences $B_{e}^{E}$ from the corresponding semantic experience bank $B_{E}$. In this context, $w$ represents a dynamic candidate related to one attribute of the instance $e\in E$, where $E\in\{\mathcal{A}+(O),ST,SG\}$ is an element of the task decomposition hierarchy $TD$. The verification of $w$ proceeds as follows: $w$ is identified as inaccurate and consequently filtered out if it does not consistently hold across all experiences within $B_{e}^{E}$ or if it conflicts with any other established dynamics. The dynamics that withstand this verification process are classified as verified dynamics, denoted as $\mathcal{W}$.
3.2 Online Strategy Learning
To effectively integrate the learned world dynamics $\mathcal{W}$ into the Crafter environments, we deploy an LLM-based agent defined by $\pi:\mathcal{S}\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{A})$. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ denotes the action space, and $\mathcal{P}$ symbolizes the probability distribution over the action space. Rather than directly mapping the world dynamics $\mathcal{W}$ and the observation $o_{t}$ of the current state $s_{t}$ to the action $a_{t}$, we address the challenge of long-horizon planning by enhancing decision-making with an online strategy learning method. This approach decomposes the planning process into three distinct tasks: sub-goal planning, sub-task planning, and action planning.
Sub-goal Planning
Given that the sub-goal sequence $SG=[sg_{1},sg_{2},\ldots]$ is derived from human demonstrations $H$ and treated as a fixed sequence, we employ a simple heuristic for sub-goal planning. When a sub-goal is completed, the current sub-goal will be updated as the first uncompleted sub-goal within $SG$.
Sub-task Planning
For a given current sub-goal $sg_{i}$, we have developed an LLM-based sub-task planner. This planner evaluates and ranks all sub-tasks $st\in ST$ based on the learned world dynamics $\mathcal{W}$, the verbalized current observation $\tilde{o}_{t}$, and the most recently planned sub-task $st_{t-1}$. The top-ranked sub-task is then set as the current sub-task $st_{t}$. To ensure precise execution, the completion of a sub-task $st$ is contingent upon meeting its specific termination condition. This condition is assessed by querying an LLM with the current verbalized observation, the verbalized observation from the time step when the sub-task began, and the termination conditions of the current sub-task.
Online Strategy Deriving
In addition to learning the basic rules of the Crafter environment (i.e., world dynamics $\mathcal{W}$), we also explore the development of advanced game-playing strategies based on these dynamics. Unlike learning the world dynamics, the strategy space is often too vast to fully explore. To address this, we propose evolving the dynamics into strategies $\mathcal{I}$ using an online learning method. This method reduces the search space by conditioning not only on the dynamics $\mathcal{W}$ but also on the verbalized current observation $\tilde{o}_{t}$ and the sub-task $st_{i}$. This targeted approach helps generate strategies that are more grounded and responsive to current game scenarios than those developed through offline methods. To facilitate this, we have designed an LLM-based Evolver that develops strategy candidates $\tilde{\mathcal{I}}$ through deductive reasoning applied to $\mathcal{W}$. Specifically, the Evolver derives strategy candidates by rules of inference, such as modus ponens. These strategy candidates $\tilde{\mathcal{I}}$ are evaluated for their validity and are ranked based on their usefulness by an LLM-based critic. Lastly, the valid and useful candidates are added to the situational strategy set $\mathcal{I}$.
Action Planning
The final action selection process is carried out in two primary steps:
  1. 1.
    Invalid Action Masking: This step involves masking all actions that are not feasible under the current game conditions, based on the verified dynamics $\mathcal{W}$ and current observation $\tilde{o}_{t}$.
  2. 2.
    Action Selection: From the set of valid actions, a specific primitive action $a$ is chosen. This selection is based on several factors: the current sub-task $st_{i}$, the verbalized current observation $\tilde{o}_{t}$, the world dynamics $\mathcal{W}$, a windowed history of previously planned actions and observations, and the strategies derived $\mathcal{I}$.
4 Experiment
In this section, we start by detailing our experimental settings including evaluation metrics and the set of baselines we are comparing against (Section 4.1). Then, we report the main results (Section 4.2). Additionally, we conduct a set of controlled experiments, we provide quantitative and qualitative analyses to obtain better understanding on DiVE (Section 4.3 and 4.4).
4.1 Environment Settings
Environment Crafter Hafner (2021) is a set of open-world survival games on $64\times 64$ grid maps, featuring a diverse array of materials such as tree, stone, and coal, as well as entities including cow, zombie, and skeleton semi-randomly spawn on the maps. The games include an achievement graph with 22 unique achievements across 7 levels. The agent observes its surroundings through a local $7\times 9$ window and is also aware of its status within the game environment.
Verbalizer
The text description from the verbalizer includes: the nearest object of each type within the accumulated observations, the objects located between these nearest objects, the objects in each direction, and the agent’s current inventory and status. We provide an example of the verbalization process in Appendix D.
Evaluation metrics
Following previous works Hafner (2021); Wu et al. (2024b), we evaluate the agent with two primary metrics: reward and score. Agents receive a $+1$ reward for each new achievement unlocked (e.g., make wood pickaxe, place furnace). Additionally, they earn a ±$\pm 0.1$ reward for every health point gained or lost. The score metric is computed by aggregating the success rates for each achievement:
$S\doteq\exp\left(\frac{1}{N}\sum_{i=1}^{N}\ln\left(1+s_{i}\right)\right)-1.$ (4)
Baseline systems
We compare DiVE with three categories of baselines. First, LLM-based methods, including SPRING Wu et al. (2024b), ELLM Du et al. (2023), and chain-of-thought (CoT) Wei et al. (2022). Second, reinforcement learning (RL) based approaches such as DreamerV3 Hafner et al. (2023), PPO Schulman et al. (2017), AD  Moon et al. (2024). Third, we consider demonstrations from human experts, which provide insights from actual human players.
Method Score Reward
Human Experts $50.5\pm 6.8\%$ $14.3\pm 2.3$
DiVE $\mathbf{35.9\pm 3.2\%}$ $\mathbf{14.5\pm 2.4}$
SPRING $8.2\pm 2.3\%$ $6.9\pm 1.8$
CoT $1.3\pm 0.3\%$ $2.5\pm 0.5$
AD $21.79\pm 1.4\%$ $12.6\pm 0.3$
ELLM N/A $6.0\pm 0.4$
DreamerV3 $14.5\pm 1.6\%$ $11.7\pm 1.9$
PPO $4.6\pm 0.3\%$ $4.2\pm 1.2$
Random $1.6\pm 0.0\%$ $2.1\pm 1.3$
Table 1: Primary table illustrating the performance of DVE in comparison to all baseline models. The symbol indicates results obtained by running the official code implementation using the same five random seed and same model.
4.2 Overall Results
Table 1 demonstrates that DiVE significantly outperforms all other baselines across two distinct evaluation metrics. Notably, DiVE exceeds the previous state-of-the-art (SOTA) LLM-based method, SPRING [^1] Results were obtained by executing the official SPRING code using the same LLM configuration (GPT-4o) and environment random seeds as DiVE, ensuring a fair comparison., by a substantial margin, achieving a 337.8% relative improvement in score and a 110.1% enhancement in reward. Additionally, DiVE also surpasses the prior SOTA RL-based approach, DreamerV3, with a 21.4% absolute improvement in score and a 2.8 absolute increase in reward. Notably, DiVE achieves rewards comparable to domain-familiar human players using just 10 human demonstrations.
4.3 Quantitative Analysis
Methods Score Reward
Component analysis
DiVE $\mathbf{35.9\pm 3.2\%}$ $\mathbf{14.5\pm 2.4}$
w/o E $21.1\pm 9.7\%$ $11.3\pm 4.3$
w/o V $9.8\pm 1.0\%$ $10.1\pm 0.7$
w/o V&E $11.5\pm 4.9\%$ $8.3\pm 3.8$
w/o D&V&E $0.9\pm 0.1\%$ $2.5\pm 1.3$
CoT $1.3\pm 0.3\%$ $2.5\pm 0.5$
CoT + D&V $3.6\pm 0.9\%$ $3.9\pm 2.3$
Dynamics from distinct sources
DiVE $\mathbf{35.9\pm 3.2\%}$ $\mathbf{14.5\pm 2.4}$
w/o D&V+S $15.7\pm 5.3\%$ $8.9\pm 5.1$
w/o D&V&E+S $12.1\pm 4.6\%$ $8.7\pm 3.0$
w/o D&V&E+H $34.2\pm 2.8\%$ $14.5\pm 0.9$
Table 2: Different components’ impact on the performance; The abbreviations D, V, and E stand for Discover, Verifier, and Evolver, respectively; S means the dynamics (manual) used in the SPRING paper; H means the human-annotated dynamics.
Ablation study
We conduct a set of ablation study to help elucidate how individual elements contribute to the overall performance of the method. As shown in the first section of Table 2, various variants of DiVE have been designed to assess the effectiveness of each component within the methodology. Notably, the significant performance disparity between DiVE and its variant lacking the Evolver component empirically underscores the Evolver’s role in effectively developing gameplay strategies based on world dynamics $\mathcal{W}$, thereby enhancing the agent’s proficiency in the Crafter environment. Similarly, the performance decline observed in the variant without the Verifier highlights the importance of formulating strategies $\mathcal{I}$ based on accurate world dynamics $\mathcal{W}$. Moreover, the additional performance reduction in the version missing both Verifier and Evolver demonstrates their orthogonal functions—precision in dynamics from the Verifier and advancement in strategies from the Evolver.
The performance of DiVE without Discoverer, Verifier, and Evolver reverts to the CoT baseline, indicating that merely decomposing the task according to the hierarchy $\mathcal{H}$ without incorporating domain knowledge $\mathcal{K}_{\text{target}}$ offers no advantage. The significant gap between CoT + D&V and DiVE w/o E further evinces that an LLM-based agent struggles in long-horizon planning tasks without task decomposition, emphasizing the necessity of the decomposition hierarchy $\mathcal{H}$.
Sources of dynamics
We also explore how DiVE performs with world dynamics derived from various sources. As shown in the second section of Table 2, DiVE significantly outperforms the variants that utilize dynamics $S^{{\dagger}}$ as used in the SPRING paper Wu et al. (2024b). This performance improvement suggests that the learned dynamics $\mathcal{W}$ are more beneficial than $S^{{\dagger}}$, possibly because $S^{{\dagger}}$ lacks some of the helpful details present in $\mathcal{W}$. We will delve deeper into this aspect in upcoming case studies. The performance difference between methods using $S^{{\dagger}}$ with and without the Evolver also underscores the importance of strategy evolution, the effectiveness of which is closely linked to the quality of the world dynamics it employs.
Besides dynamics learned from human demonstrations and the manual dynamics used in SPRING, we have also investigated a third source: human-annotated dynamics. The findings reveal that DiVE performs comparably to the variants that use human-annotated dynamics, demonstrating the robustness and effectiveness of DiVE’s approach to dynamic learning.
Figure 3: Recall of learned dynamics over discovery steps with mean and standard deviation.
Quality of learned dynamics
As previously mentioned, directly quantifying the desired properties is challenging since we cannot precisely measure how much domain-relevant information $\mathcal{K}_{\text{rel}}$ an LLM contains or determine the exact amount of knowledge $\mathcal{K}_{\text{target}}$ required by a domain. However, by using human-annotated dynamics $H^{{\dagger}}$ as a silver standard for $\mathcal{K}_{\text{target}}$ , we can estimate the precision and recall of the learned dynamics $\mathcal{W}$. This approach allows us to effectively track and evaluate the progress of the LLM-based dynamic learning. Specifically, we calculate recall as $R=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|H^{{\dagger}}|}$ and precision as $P=\frac{|\mathcal{W}\cap H^{{\dagger}}|}{|\mathcal{W}|}$. As illustrated in Figure 3, both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$ demonstrate an increase in recall with the progression of discovery steps, indicating that the richness of the learned dynamics enhances as more discovery steps are taken. Furthermore, the gap in recall between $\tilde{\mathcal{W}}$ and $\mathcal{W}$ narrows as the number of discovery steps increases, suggesting that the Verifier effectively filters out "noisy" dynamic candidates while retaining those that generalize across different trajectory segments.
To further investigate whether the Verifier retains correct world dynamic candidates $\tilde{\mathcal{W}}{+}$ while filtering out the inaccurate ones $\tilde{\mathcal{W}}{-}$, we examine the precision of both the discovered dynamics $\tilde{\mathcal{W}}$ and the verified dynamics $\mathcal{W}$. As shown in Figure 4, the precision of the verified dynamics consistently and significantly surpasses that of the discovered dynamics, demonstrating that the Verifier effectively identifies and eliminates inaccurate dynamic candidates. This validates the Verifier’s role in enhancing the reliability of the dynamics used for decision-making processes.
Figure 4: Precision of learned dynamics before and after dynamic verification.
4.4 Qualitative Analysis
The correctness of the learned and verified dynamics is categorized as either correct or as mistakes due to confounders, in-domain hallucinations, or out-of-domain hallucinations. As illustrated in Table 3, in the example of defeating a zombie, the mistake attributed to confounders is primarily due to the simultaneous increase in health points at the step where the zombie is defeated. In this case, the Discoverer categorizes this increase as a result of defeating the zombie. For the in-domain hallucination example, it is impossible to increase the wood while defeating the zombie, though the wood may appear in the observation. Lastly, the out-domain hallucination is responsible for discovering dynamics that contain nonexistent objects in the observation or even the Crafter environment.
Correctness Outcome
Correct ($\checkmark$) none
Confounder (✗) 1 health
In-domain Hallucination (✗) 1 wood
Out-domain Hallucination (✗) 1 bone
Table 3: The dynamics underlying the outcome of defeating zombie
Compared to the dynamics from SPRING in Table 4 , we found that DiVE’s dynamics are not only more precise but also more informed. For example, regarding the dynamics of placing a stone, SPRING only found that placing a stone requires stones; however, DiVE found that placing a stone requires exactly one stone and the precise facing condition for successful placement. Furthermore, based on this information, the Evolver can derive advanced dynamics for placing a stone, such as its potential to act as a barrier between the agent and dangerous creatures.
Methods               Dynamics
SPRING Place stone requires stones
D&V Place stone requires 1 stone and faces paths, grass, sand, water, and lava
D&V&E Place stone to block zombies and skeletons, preventing them from reaching the player
Table 4: Comparing SPRING and DiVE on place stone’s dynamics
5 Related Work
Embodied Agent   Reinforcement learning (RL) has long been an effective method for developing embodied agents through trial and error Moon et al. (2024); Hafner et al. (2023); Schulman et al. (2017); Jiang et al. (2022); Hafner et al. (2020); Hessel et al. (2018); Sekar et al. (2020); Weir et al. (2022). With recent advancements in LLMs Achiam et al. (2023); Team et al. (2023); Brown et al. (2020); Touvron et al. (2023); Anil et al. (2023), their reasoning and planning capabilities have emerged as an important component for the embodied agent Wang et al. (2023a, b); Ma et al. (2023); Liang et al. (2023); Brohan et al. (2023); Nottingham et al. (2023); Silver et al. (2024); Driess et al. (2023); Shi et al. (2024); Wu et al. (2024b).
Discovering and Verifying   LLMs can uncover knowledge by inducing and deducing rules for reasoning tasks Zhu et al. (2023). Additionally, they can directly discover the underlying domain knowledge from previous trajectories and interactive experience Colas et al. (2023); Majumder et al. (2023); Fu et al. (2024a, b); Zhong et al. (2024); Zhao et al. (2024).The discovered knowledge within these methods is not hierarchically structured and mainly focuses on discovering short-horizon span knowledge. Meanwhile, due to the tendency for hallucination Zhang et al. (2023b); Dziri et al. (2022b), the knowledge discovered should be carefully verified before being used directly.
Evolving   By reflecting on past trajectories, LLMs can enhance their decision-making processShinn et al. (2024). Leveraging this capability, works such as Wang et al. (2023a); Stengel-Eskin et al. (2024); Zhang et al. (2023a); Nottingham et al. (2024) evolve new and advanced skills based on pre-defined skills. However, previous methods often assume that LLMs already possess comprehensive domain knowledge and can directly develop new insights. Additionally, they often requrie pre-defined skills in either code format or RL polices.
6 Conclusion
In this work, we present DiVE, a framework designed for long-horizon interactive planning tasks. Specifically, given a handful of demonstrations, a Discoverer can extract useful world dynamics such as an action’s preconditions and outcomes. A Verifier subsequently filters out inaccurate dynamic candidates. In an online setting, conditioned on the offline-learned knowledge, an Evolver learns situational strategies through interaction. DiVE bridges the knowledge gap between a pre-trained LLM and the target domain by autonomously learning hierarchical world dynamics and evolving contextual knowledge. Extensive experiments demonstrate its effectiveness. This work may pave the way for developing LLM-based frameworks to achieve the mutual enhancement of dynamics learning and policy learning without extensive human annotation.
Limitations
Experimenting with DiVE in a stochastic yet simple environment may not adequately represent all the potential real-world scenarios that an embodied agent might encounter upon deployment. Additionally, DiVE requires collecting human expert trajectories, which can be either impossible or resource-intensive in certain situations. Finally, obtaining perfect object information might not be feasible in some environments.
Ethical Concerns
We do not anticipate any immediate ethical or societal impact arising from our work. In this study, we aimed to bridge the knowledge gap of LLMs towards the target domain. However, despite our efforts, using LLMs as the backbone, DiVE might still hallucinate due to the inherent tendency of LLMs to hallucinate.
Appendix
Appendix A Dynamics Example from DiVE
grass
grass can be found near [’tree’, water’, path’], but it is not associated with [’diamond’, coal’, iron’]
You can walk directly through grass.
grass can only be used for: [’collect_sapling’, eat_plant’] ,
coal
coal can be found near [’stone’, iron’, diamond’], but it is not associated with [’grass’, cow’, skeleton’]
You cannot walk directly through coal.
coal turn into path after collect_coal
coal can only be used for: [’make_iron_pickaxe’, make_iron_sword’, collect_coal’]
coal can be collected by [’wood_pickaxe’] ,
Appendix B Trajectory Example
subgoal
Collect 4 units of wood and maintain a
healthy level of health, food, drink,
and energy.
subtask
collect_wood": {
"General Plan": {
"step_1": "Locate a tree, ensuring tree is within immediate distance.",
"step_2": "Move closer to the tree if its not within immediate distance. This may involve navigating around obstacles such as grass, stone, or paths.",
"step_3": "Face the tree to meet the precondition of the action.",
"step_4": "Execute the collect_wood action to gather wood from the tree."
"termination_condition": "The inventorys wood amount increases by 1, or if health, food, drink, or energy levels drop critically low, or a threat (such as skeletons or zombies) is detected nearby."
state description
It is daytimeState description:
- East: immediate (grass); nearby (unexplored_area);
- North: immediate (grass); nearby (unexplored_area);
- North East: immediate (grass); nearby (unexplored_area);
- North West: immediate (grass); nearby (unexplored_area);
- South: immediate (grass); nearby (unexplored_area);
- South East: immediate (grass); nearby (cow, unexplored_area);
- South West: immediate (grass); nearby (unexplored_area);
- West: immediate (grass); nearby (unexplored_area);
Closest:
- grass: North 1 blocks away (immediate) (objects in between: None)
- cow: South East 5 blocks away (nearby) (objects in between: {’grass’})
- Facing grass on the south.
Your status:
- health: 9/9
- food: 9/9
- drink: 9/9
- energy: 9/9
You have nothing in your inventory.
evolve dynamics
Expand the exploration area to discover new surroundings that could potentially include a tree.
action
move_north
Appendix C Precision changes over time
Figure 5: Precision of verified dynamics over verified steps
As shown in Figure 5, the precision of the verified dynamics $\mathcal{W}$ increases during the initial step of verification. However, as the verification process progresses, further increases in precision are not observed. This suggests that the first step effectively filters out all the unreliable dynamics.
Appendix D Verbalizer
Figure 6: The observation within the Crafter environment
Verbalized Observation
It is daytime
State description:
- East: immediate (grass); nearby (unexplored_area);
- North: immediate (grass); nearby (unexplored_area);
- North East: immediate (grass); nearby (unexplored_area);
- North West: immediate (grass); nearby (tree, unexplored_area);
- South: immediate (grass); nearby (unexplored_area);
- South East: immediate (grass); nearby (unexplored_area, tree);
- South West: immediate (grass); nearby (tree, unexplored_area);
- West: immediate (grass); nearby (tree, unexplored_area);
Closest:
- grass: North 1 blocks away (immediate) (objects in between: None)
- tree: West 4 blocks away (nearby) (objects in between: {’grass’})
- Facing grass on the south.
Your status:
- health: 9/9
- food: 9/9
- drink: 9/9
- energy: 9/9
You have nothing in your inventory.
Appendix E Crafter’s Tech Tree
Figure 7: The tech tree within the Crafter environment
Appendix F Dynamic discovering and verification algorithm
Input: Verbalized human demonstrations $\tilde{H}$, current element $E\in TD$ for discovering, dynamics of $E$’s parents $\{\mathcal{W}_{V}\}_{V\in\text{Parents}(E)}$
Constant: Discovery steps $s$, action space $\mathcal{A}$, object set $O$, semantic trajectory transformation function $T^{E}$, dynamic distillation function $D^{E}$, semantic sampling function $S_{E}$.
Result: $\tilde{\mathcal{W}}_{E}$
$\tilde{\mathcal{W}}_{E}\leftarrow\{\}$;
$B^{E}\leftarrow T^{E}(\tilde{H})$ ;
  /* $B^{E}$ is the semantic experience bank. */
for $k\leftarrow 1$ to $s$ do
       for $e\in E$ do
             for $x\in n.attributes$ do
                   $B_{e}^{E}\leftarrow S_{E}(n,B^{E})$;
                   $y\leftarrow D^{E}(x,B_{e}^{E},\{\mathcal{W}_{V}\})$;
                   $\tilde{\mathcal{W}}_{E}\leftarrow\tilde{\mathcal{W}}_{E}\cup\{y\}$
             end for
       end for
end for
Algorithm 1 Dynamics discovering
Input: Dynamic candidate set $\tilde{\mathcal{W}}_{E}$, current level $E\in TD$ for verification, semantic experience bank $B^{E}$ concerning $E$, dynamics of $E$’s parents $\{\mathcal{W}_{V}\}_{V\in\mathbf{Parents}(E)}$
Constant: Dynamic verifying function $V^{E}$, conflict detection function $F^{E}$, semantic sampling function $S_{E}$
Result: $\mathcal{W}_{E}$
$\mathcal{W}_{E}\leftarrow\{\}$;
$\bar{\mathcal{W}}_{E}\leftarrow\{\}$;
for $w\in\tilde{\mathcal{W}}_{E}$ do
       $B^{E}_{w}\leftarrow S_{E}(w,B^{N})$;
       if $V^{E}(w,B_{w}^{E})$ is True then
             $\bar{\mathcal{W}}\leftarrow\bar{\mathcal{W}}_{E}\cup\{w\}$
       end if
end for
for $w\in\bar{\mathcal{W}}_{E}$ do
       if $F^{E}\left(w,\bar{\mathcal{W}}_{E},\{\mathcal{W}_{V}\}\right)$ is False then
             $\mathcal{W}_{E}\leftarrow\mathcal{W}_{E}\cup\{w\}$
       end if
end for
Algorithm 2 Dynamics verfication
Appendix G Prompts
G.1 Grounding
G.1.1 Subtask Selection
Given the following details:
- Subtask description: {self._transition[’subtask’]},
- Current observation: {self._transition[’state_description’]},
- Initial observation: {self._transition[’initial_state_description’]},
- Previous executed actions: {self._transition[’previous_actions’][-3:]}
you are asked to decide whether the subtask should be terminated or not.
For deciding whether to terminate the subtask, consider:
- The previous action, provided it was executed successfully.
- The difference between the initial and current observations, including the inventory changes.
The subtask should be terminated, and the output should be True only if any of its termination conditions are met.
Otherwise, if none of the termination conditions are met, the subtask should continue running, and the output should be False’.
Justify whether the termination conditions are met or not first, and then provide the termination decision.
Output in this format: {output_format}.
G.1.2 Online Evolve
Given the following details:
- Primitive dynamics: {env_dynamics}
- Current subtask: {transition[’subtask’]}
- Current observation: {transition[’state_description’]}
You are asked to identify the difficulties in completing the current subtask and provide 3 primitive or evolved dynamics to solve each of these difficulties.
Instructions for identifying difficulties:
- List the objects required to complete the subtask, specify their locations, and explain where to find them if they are not in the current observation.
- Outline all possible obstacles that may be encountered along the way.
Instructions for evolving advanced dynamics:
- Do not introduce any new objects that are not part of the primitive dynamics.
- The evolved dynamics should not contradict the primitive dynamics.
- If a difficulty cannot be resolved by existing dynamics, evolve new and advanced dynamics by combining only existing dynamics using deductive reasoning.
Instructions for providing deductive reasoning steps:
- For each evolved dynamics, provide the used primitive dynamics.
- For the deductive reasoning steps, provide the steps to combine the primitive dynamics to evolve the advanced dynamics and the rule of inference used (Modus Ponens, Modus Tollens, ......)
Last, for each situation and dynamics should be general and do not contain details about specific locations about the objects.
Output in the following format:{output_format} and leaves None for deductive_reasoning_steps if the dynamics are primitive.
G.2 Critics for Evolved Dynamics
Given the following details:
- Primitive dynamics: {env_dynamics}
- Evolved dynamics: {transition[’reformatted_dynamics’]}
- Current subtask: {transition[’subtask’]}
You are asked to examine the validity of the evolved dynamics and provide feedback.
Instructions for examining the validity of the evolved dynamics:
- The evolved dynamics should only be a combination of existing dynamics using deductive reasoning and should not introduce new dynamics. Output True if the evolved dynamics introduce new dynamics; otherwise, output False’.
- The evolved dynamics should not introduce any new objects that are not mentioned in the primitive dynamics. Output True if the evolved dynamics introduce new objects; otherwise, output False’.
- Each deductive reasoning step should not contradict any of the primitive dynamics. Output True if the evolved dynamics contradict the primitive dynamics; otherwise, output False’.
Instructions for examing the usefulness of the evolved dynamics:
- The difficulties should be directly related to the current subtask.
- The evolved dynamics should be useful in solving the difficulties identified in the subtask.
- Evluating the usefulness of the evolved dynamics on a scale of 1 to 5, where 5 is the most useful and 1 is the least useful.
Last, output the validity of the evolved dynamics in the following format: {output_format}.
G.2.1 Action Selection
Given the following details:
- Current observation: {self._transition[’state_description’]}
- Current subtasks description: {self._transition[’subtask’]}
- Previous actions: {self._transition[’previous_actions’][-3:]}
- Primitive dynamics: {self._transition[’primitive_dynamics’]}
You are asked to:
- identify the objects related to the current subtask and provide their locations and dynamics.
- select the top 3 actions that contributes to the subtask by either moving closer to the object or interacting with the object; and provide all the objects and dynamics directly related with each action.
- based on each actions related objects, provide the rationale and detailed consequences of executing each action on the objects.
- select the best action to execute next and provide the justification for your choice.
Lastly, select the action only from the available actions: {self._transition[’available_actions’]}; {feedback}.
Note: Avoid unnecessary crafting and placement if the items are within reachable distance.
Please format your response in the following format: {action_format}
G.3 Discover Action Dynamics
G.3.1 Materials used attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the materials used in the inventory for executing this action.
To discover the materials used in the inventory, consider the following aspects:
- List all the materials used not gained for each state transition and
their corresponding quantities;
- List all the common materials used in all state transitions and their corresponding
quantities.
If no materials are used in common, simply output as "None".
materials are: wood, stone, coal, iron, diamond, sapling.
Output as many pre-conditions as you can think of in the following format: {output_format}
G.3.2 Immediate objects attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the objects within immediate distance required for executing this action.
To discover the objects within immediate distance, consider the following aspects:
- List all the objects within immediate distance for each state transition before executing the action, which you can only choose from the object list.
- Identify and list the objects that are present in all state transitions within the immediate distance before executing the action, which you can only choose from the object list.
- The common object **must** be present in all state transitions immediate objects.
If no objects within immediate distance are in common, simply output as "None".
Object list: [coal, cow, diamond, furnace, iron, lava, skeleton, stone, table, tree, water, zombie, plant]
Output as many pre-conditions as you can think of in the following format: {output_format}
G.3.3 Facing object attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the facing object required for executing this action.
To discover the objects within immediate distance, consider the following aspects:
- List the facing object for each state transition before this action executed, not the direction.
- List the union of all the facing object across all the state transitions before this action executed.
- The union of all the facing object must be present in at least one state transitions facing object.
Object list: [coal, cow, diamond, furnace, iron, lava, skeleton, stone, table, tree, water, zombie, plant]
Output as many pre-conditions as you can think of in the following format: {output_format}
G.3.4 Inventory tool attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the the inventory tool required for executing this action.
To discover the objects within immediate distance, consider the following aspects:
- List all the tools in the inventory for each state transition.
- List the tools that are common across all the state transitions.
- The common tools must be present in all state transitions tools; if no tools are required, simply output as "None".
- List the most advanced tool required for executing the action within the common tools.
Tool list: [None, wood pickaxe, stone pickaxe, iron pickaxe, wood sword, stone sword, iron sword].
Output as many pre-conditions as you can think of in the following format: {output_format}
G.3.5 Outcome attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the inventory and status increase after executing this action.
To discover the increase about the inventory and status, consider the following aspects:
- List all the increase of the inventory and status for each state transition after executing the action.
- List the increase of the inventory and status that are common across all the state transitions after executing the action.
- The common increase must be present in all state transitions increase; if no inventory and status can be found, simply output as "None";
Output as many pre-conditions as you can think of in the following format: {output_format}
G.3.6 Facing object change attribute
In the Crafter environment, the world dynamics are unique.
You are tasked with discovering the unique dynamics of actions within this environment.
Given the state transitions described as follows: {partial_description}, identify the facing object changes for executing this action.
To discover the change about the facing object, consider the following aspects:
- List all the change about the facing object after executing the action for each state transition.
- List the common change about the facing object across all the state transitions.
- The common change must be present in all state transitions facing object change; if no change can be found, simply output as "None".
inventory: [wood, stone, coal, iron, diamond, sapling, wood_pickaxe, stone_pickaxe, iron_pickaxe, wood_sword, stone_sword, iron_sword]
status: [health, food, drink, energy], the maximum value of status is 9 and you can use increased_to_9 to represent the status increase to the maximum value.
Output as many pre-conditions as you can think of in the following format: {output_format}
G.4 Verify Action Dynamics
G.4.1 Materials used attribute
In the crafter environment, you have discovered unique dynamics and need to verify the pre-conditions for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the inventory materials used pre-conditions for the action ’{action}’ based on the discovered dynamics.
Precondition: You need to use {precondition} to execute the action.
For the verification, you need to consider the followings:
- List all the inventory materials that are consumed not gained for each state transition and their corresponding quantity,
- List the common inventory materials used across all the state transitions.
- Determine if the inventory materials used mentioned in the pre-condition are within the common inventory materials used.
- If it is within the common inventory materials used, then it is valid; otherwise, it is invalid.
Finally, if there are more advanced tools exist, then it is invalid; otherwise, it is valid.
materials are: wood, stone, coal, iron, diamond, sapling.
Output in this format: {output_format}
G.4.2 Immediate objects attributes
In the crafter environment, you have discovered unique dynamics and need to verify the pre-conditions for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the immediate objects pre-conditions for the action ’{action}’ based on the discovered dynamics.
Precondition: You only need {precondition} within immediate distance.
For the verification, you need to consider the followings:
- List all the objects within immediate distance before executing the action for each state transition only from the object list.
- Identify and list the objects that are present in all state transitions within the immediate distance before executing the action as common immediate objects, only from the object list.
- The common object **must** be present in all state transitions immediate objects.
- Determine if there are objects mentioned in the pre-condition that are not in the common immediate objects.
- If there are objects mentioned in the pre-condition that are not in the common immediate objects, then it is invalid; otherwise, it is valid.
Object list: [coal, cow, diamond, furnace, iron, lava, skeleton, stone, table, tree, water, zombie, plant]
Output in this format: {output_format}
G.4.3 Facing object attribute
In the crafter environment, you have discovered unique dynamics and need to verify the pre-conditions for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the facing object pre-conditions for the action ’{action}’ based on the discovered dynamics.
Precondition: You need to face {precondition} before execute the action.
For the verification, you need to consider the followings:
- List the facing object for each state transition before this action executed.
- List the union of all the facing object across all the state transitions before this action executed.
- Determine if the facing object mentioned in the pre-condition is within the union of all the facing object.
- If the facing object mentioned in the pre-condition is within the union of all the facing objects, then it is valid; otherwise, it is invalid.
Finally, if there are more advanced tools exist, then it is invalid; otherwise, it is valid.
Object list: [coal, cow, diamond, furnace, iron, lava, skeleton, stone, table, tree, water, zombie, plant]
Output in this format: {output_format}
G.4.4 Inventory tool attribute
In the crafter environment, you have discovered unique dynamics and need to verify the pre-conditions for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the inventory pre-conditions for the action ’{action}’ based on the discovered dynamics.
Precondition: You only need {preconditions} to execute the action.
For the verification, you need to consider the followings:
List all the tools in the inventory for each state transition.
- List the tools that are common across all the state transitions.
- The common tools must be present in all state transitions tools; if no tools are required, simply output as "None".
- If the tools mentioned in the pre-condition are within the common tools, then it is valid; otherwise, it is invalid.
Tool list: [None, wood pickaxe, stone pickaxe, iron pickaxe, wood sword, stone sword, iron sword].
Output in this format: {output_format}
G.4.5 Outcome attribute
In the crafter environment, you have discovered unique dynamics and need to verify the outcome for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the inventory and status increase after executing this action. ’{action}’ based on the discovered dynamics.
predicted_increases: You only increase {precondition} in the inventory and status after executing the action.
For the verification, you need to consider the followings:
- List all the increases of the inventory and status for each state transition after executing the action as the increases.
- List the increases of the inventory and status that are common across all the state transitions after executing the action as the common increases.
- The common increases must be present in all state transitions increases; if no inventory and status increases can be found, simply output as "None".
- Determine if there are increases mentioned in the predicted_increases that are not in the common increases.
- If there are increases mentioned in the predicted_increases that are not in the common increases, then it is invalid; otherwise, it is valid.
Output in this format: {output_format}
inventory: [wood, stone, coal, iron, diamond, sapling, wood_pickaxe, stone_pickaxe, iron_pickaxe, wood_sword, stone_sword, iron_sword]
status: [health, food, drink, energy], the maximum value of status is 9 and you can use increased_to_9 to represent the status increase to the maximum value from the non-maximum value.
G.4.6 Facing object change attribute
In the crafter environment, you have discovered unique dynamics and need to verify the outcome for the action ’{action}’.
Given the state transitions before and after taking the action ’{action}’, described as follows: {sampled_descriptions}
You are asked to verify the facing object changes after executing this action. ’{action}’ based on the discovered dynamics.
predicted_changes: The facing object changes is {precondition} after executing the action.
For the verification, you need to consider the followings:
- List all the change about the facing object after executing the action for each state transition.
- List the common change about the facing object across all the state transitions.
- The common change must be present in all state transitions facing object change; if no change can be found, simply output as "None".
- Determine if the predicted changes are within the common changes.
- If the predicted changes are within the common changes, then it is valid; otherwise, it is invalid.
G.5 Discover Objects Dynamics
Given randomly selected frames featuring ’{obj}’, please complete the following:
1. Identify which objects are most closely related to ’{obj}’.
- Rate their relevance using these terms: Very related’, Not related’.
- Here are the selected frames concerning ’{obj}’: {description}
2. Identify the common time relationship for this object.
3. Format your response as follows: {object_relationship_format}
Only consider the relationship with the following objects: [’grass’, coal’, cow’, diamond’, iron’, lava’, skeleton’, stone’, tree’, water’, zombie’, plant’, path’, sand’, plant-ripe’]
For the relationship definitions:
- Very related: They can always be found together.
- Not related: They cannot be found together.
G.6 Verify Objects Dynamics
Given randomly selected frames featuring ’{obj}’ described as: {description}, please complete the following tasks:
1. Identify the objects and time most closely related to ’{obj}’ in the provided description.
- Rate their relevance using the following terms: Very related, Not related’.
2. Based on the discovered dynamics {dynamics}, verify the relationships between the objects and time.
3. Output all the valid relationships within the discovered dynamics that match all the relevance levels of the objects in the provided frames only about the object ’{obj}’.
Consider relationships with the following objects only: [’grass’, coal’, cow’, diamond’, furnace’, iron’, lava’, skeleton’, stone’, table’, tree’, water’, zombie’, plant’, path’, sand’, plant-ripe’].
Relationship definitions:
- Very related: They are always found together.
- Not related: They are never found together.
Output in the following format: {object_relationship_format}
G.7 Discover Sub-task Dynamics
In the Crafter environment, the world dynamics are complex and there are many subtasks to be completed.
Here are the dynamics about the core action {action}:
The precondition:
- it need to face: {self.facing_object_preconditions[action]}
- it need to have {self.immediate_object_preconditions[action]} within immediate distance
- it need to have {self.inventory_materials_precondition[action]} in the inventory.
- it need to have {self.inventory_tool_precondition[action]} in the inventory.
The outcome:
- its facing object changes to {self.facing_object_change[action]}
- its inventory and status outcome increases on {self.inventory_outcome[action]}
Note: None indicates that there are no specific preconditions or outcomes for the corresponding elements of the dynamics."
Given the subtask "{subtask}", I want you to write the step-plan for completing this subtask.
First, read and understand all the provided world dynamics.
Then, locate the world dynamics that are relevant to the subtask.
Next, write the subtasks requirements, the step-plan for completing the subtask based on the world dynamics and the outcomes in this format {subtask_plan_format}
-steps: the general steps required to complete the subtask.
-termination_condition: when should this subtask be terminated; you should consider the status, potential danger and the outcome of the action
Finally, only output the plan for the subtask.
Here are a few examples: {examples}
Here are the observations on how to completing the subtask: {partial_description}
G.8 Verify Sub-task Dynamics
In the Crafter environment, the world dynamics are complex and there are many subtasks to be completed.
Here are the dynamics about the core action {action}:
The precondition:
- it need to face: {self.facing_object_preconditions[action]}
- it need to have {self.immediate_object_preconditions[action]} within immediate distance
- it need to have {self.inventory_materials_precondition[action]} in the inventory.
- it need to have {self.inventory_tool_precondition[action]} in the inventory.
The outcome:
- its facing object changes to {self.facing_object_change[action]}
- its inventory and status outcome increases on {self.inventory_outcome[action]}
Note: None indicates that there are no specific preconditions or outcomes for the corresponding elements of the dynamics."
Given the subtask "{subtask}", I want you to verify the plan for completing this subtask {discovered_plan}.
First, read and understand all the provided world dynamics.
Then, locate the world dynamics that are relevant to the subtask.
Next, examine the provided plans requirements, the step-plan and the termination conditions.
Lastly, if it can be applied across different examples, then it is a valid plan.
Here are the observations on how to completing the subtask: {partial_description}
Output within this format{subtask_verification_format}
G.9 Discover Sub-goal Dynamics
Given the following considerations:
- The games final goal: collect diamonds and survive.
- All the subtasks: {[’collect_coal’, collect_diamond’, collect_drink’, collect_iron’, collect_sapling’, collect_stone’, collect_wood’, defeat_skeleton’, defeat_zombie’, eat_cow’, eat_plant’, make_iron_pickaxe’, make_iron_sword’, make_stone_pickaxe’, make_stone_sword’, make_wood_pickaxe’, make_wood_sword’, move’, place_furnace’, place_plant’, place_stone’, place_table’, sleep’]}
- Environment dynamics: {env_dynamics}
- Human player completed subtasks in chronological order: {human_demo}
Note: action_x_y means this action is performed consecutively for y times
Your task is to generate a plan by reordering the subtasks to help the player achieve the final game goal with different subgoals.
For plan generation, consider:
- List all the subtasks that contributes to the final goal from the human players trajectory.
- Re-order the subtasks that fits the games dynamics.
- For each subtask, state it as a subgoal with a description and the exact repetition of the subtask if needed.
Lastly, output the plan in this format: {output_format};
For example: {output_example}
G.10 Verify Sub-goal Dynamics
Given the following considerations:
- The games final goal: collect diamonds and survive.
- All the subtasks: {[’collect_coal’, collect_diamond’, collect_drink’, collect_iron’, collect_sapling’, collect_stone’, collect_wood’, defeat_skeleton’, defeat_zombie’, eat_cow’, eat_plant’, make_iron_pickaxe’, make_iron_sword’, make_stone_pickaxe’, make_stone_sword’, make_wood_pickaxe’, make_wood_sword’, move’, place_furnace’, place_plant’, place_stone’, place_table’, sleep’]}
- Environment dynamics: {env_dynamics}
- Human player completed subtasks in chronological order: {human_demo}
Note: action_x_y means this action is performed consecutively for y times
- Discovered plan: {discovered_plan}
Your task is to verify the provided plan.
For plan verification, consider:
- List all the subtasks that contributes to the final goal from the human players trajectory.
- Re-order the subtasks that fits the games dynamics.
- If the provided plan can be applicable to the different trajectories, then it is valid.
Lastly, output this format: {output_format};